OpenStudy (help!!!!):

2. If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together? A. 2 hours and 24 minutes B. 3 hours and 12 minutes C. 3 hours and 44 minutes D. 4 hours and 10 minutes E. 4 hours and 33 minutes

OpenStudy (help!!!!):

this is how far i am 1/4+1/6=5/12 then what is next after this?

jimthompson5910 (jim_thompson5910):

1/4 + 1/6 = 1/t 5/12 = 1/t 5t = 12 keep going to solve for t to get your final answer

OpenStudy (help!!!!):

how did you know it was 1/t?

jimthompson5910 (jim_thompson5910):

because if they get the job done together in t hours, then their combined rate is 1/t jobs per hour

OpenStudy (help!!!!):

how could i remeber that somehow?

jimthompson5910 (jim_thompson5910):

you can think of it as rate * time = (amount of work done) and you can think of "amount of work done" to be the value of 1 (to signify you have done 1 full completed job)

jimthompson5910 (jim_thompson5910):

then you solve for the rate to get rate = (amount of work done)/(time) r = 1/t

OpenStudy (help!!!!):

thank guys for coming you guys are alway there when i need help thxss

jimthompson5910 (jim_thompson5910):

np

OpenStudy (help!!!!):

oh yeah how did you guys get so many points?

jimthompson5910 (jim_thompson5910):

just by answering questions

OpenStudy (help!!!!):

how mnay years? or months?

jimthompson5910 (jim_thompson5910):

idk lost count lol, but a while though idk if it's been years though

OpenStudy (help!!!!):

lol